Sol Summary – December 27th

Sol: 3
Summary Title: Sol Three: Bivalves for me!
Author’s Name: Adriana Brown
Mission Status: Nominal
Sol Activity Status:
Crew 289 awoke at 7:30 am and ate a breakfast of berries and biscuits by Adriana (CO). The crew then engaged in morning stretches to limber up for a day of research and fieldwork. At 10:00 AM, Adriana led a briefing for the EVA crew, Eshaana (GEO), Gabe (HSO), Riya (GHO), and Adriana, on what they would be looking for in the field in terms of specimens, formations, documentation, and the sample collection plan. The two main objectives of the EVA were to collect bivalve fossils from the Tununk Shale and locate bentonite beds in the area. The EVA crew departed at 11:00 AM, quickly awed by the colorful strata, views of the sprawling Martian landscape, and ancient shells scattered about. The EVA crew met their objectives, collecting 36 oyster fossils from two sites on Hab Ridge, noting the coordinates, and also sampled the sediment matrices of the two sites and a bentonite bed. Riya gathered LiDAR scans of Hab Ridge that will be processed this evening.
While on the EVA, the Hab crew made a delicious cheesy broccoli soup (Adi) with homemade bread (Sara), and with some remnant spaghetti noodles (Nathan), as well as worked on robots and processed survey data. Once the EVA crew returned, the team had a group lunch then dispersed to work on research. Eshaana troubleshooted her mini farm sensor and labeled its plants, Adi worked on his robot son, E.L.F. and Nathan soldered a battery pack for his air quality sensors.
This evening, Gabe is gearing up to compile his photometry measurements for his variable star and set a baseline for his dust experiment (aka cleaning Hab surfaces!). Adi is brainstorming mounts so the drone can take LiDAR scans, and Adriana will check on her samples drying in the Science Dome. Dinner will be provided by Nathan and Gabe which will feature pasta and chickpeas with a side of report writing.
Look Ahead Plan: Geologic EVA #5 to Sea of Shells
Anomalies in work: None
Weather: Cold, clear, and sunny
Crew Physical Status: Nominal
EVA: geologic EVA #4 to Hab Ridge
Reports to be filed: Journalist Report, Greenhab Report, Operations Report, Sol Summary, Astronomy Report, Photos, Mission Summary Report
Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site